Suppose (X,Y,Z)^T has exponential distribution with mean vector (1,2,3)^T . Find the correlation matrix of (X,Y,Z)^T .


1
Expert's answer
2021-10-10T18:25:24-0400

The probability density function (pdf) of an exponential distribution is

"f(x,\lambda)=\begin{cases}
 \lambda e^{-\lambda x} &x\ge0 \\
 0&x<0
\end{cases}"



correlation coefficients:


between X and Y:

"r_{XY}=\frac{cov(X,Y)}{\sigma_X \sigma_Y}"

for exponential distribution: "\sigma_X=E(X)"

"cov(X,Y)=E(XY)-E(X)E(Y)=E(XY)-2"

"r_{XY}=\frac{E(XY)-2}{2}"


between y and z:

"r_{YZ}=\frac{cov(Y,Z)}{\sigma_Y \sigma_Z}"

for exponential distribution: "\sigma_Y=E(Y),\ \sigma_Z=E(Z)"

"cov(Y,Z)=E(YZ)-E(Y)E(Z)=E(YZ)-6"

"r_{YZ}=\frac{E(YZ)-6}{6}"


between y and z:

"r_{XZ}=\frac{cov(X,Z)}{\sigma_X \sigma_Z}"

"cov(X,Z)=E(XZ)-E(X)E(Z)=E(XZ)-3"

"r_{XZ}=\frac{E(XZ)-3}{3}"


Correlation matrix: